# Exploring Knowledge Boundaries in Large Language Models for Retrieval Judgment

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly recognized for their practical applications. However, these models often encounter challenges in dynamically changing knowledge, as well as in managing unknown static knowledge. Retrieval-Augmented Generation (RAG) tackles this challenge and has shown a significant impact on LLMs. Actually, we find that the impact of RAG on the question answering capabilities of LLMs can be categorized into three groups: beneficial, neutral, and harmful. By minimizing retrieval requests that yield neutral or harmful results, we can effectively reduce both time and computational costs, while also improving the overall performance of LLMs. This insight motivates us to differentiate between types of questions using certain metrics as indicators, to decrease the retrieval ratio without compromising performance. In our work, we propose a method that is able to identify different types of questions from this view by training a Knowledge Boundary Model (KBM). Experiments conducted on 11 English and Chinese datasets illustrate that the KBM effectively delineates the knowledge boundary, significantly decreasing the proportion of retrievals required for optimal end-to-end performance. Specifically, we evaluate the effectiveness of KBM in three complex scenarios: dynamic knowledge, long-tail static knowledge, and multi-hop problems, as well as its functionality as an external LLM plug-in.
@@ 1 Introduction
As Large Language Models (LLMs) evolve, their real-world applications expand, yet they often struggle with dynamically changing and unknown static knowledge, leading to inaccuracies or hallucinations (Rawte et al., 2023). Retrieval-Augmented Generation (RAG) effectively addresses these challenges by retrieving relevant external information in real time, enhancing LLMs’ ability to provide accurate responses. While RAG can significantly boost performance, it also incurs costs, such as increased retrieval requests and longer response times. This raises a crucial question: when is retrieval truly necessary?
@@
Previous studies on the necessity of RAG for LLMs can be categorized into two main approaches. The first focuses on the query itself, with methods like SELF-RAG (Asai et al., 2023) instructing models such as GPT-4 (Achiam et al., 2023) to assess whether retrieving external documents (e.g., Wikipedia) can produce better responses. Although this approach based on query instructing can identify questions that require real-time information, it remains model-agnostic and struggles to determine whether an LLM has mastered specific knowledge. The second approach evaluates both questions and model responses to determine if an LLM can answer a question, generating data by sampling multiple model responses and using manual labels for evaluation. However, this method is labor-intensive and relies heavily on manual labeling, which can create biases and lead to increased training costs. Recent advancements in adaptive retrieval (Jeong et al., 2024) highlight the importance of tailoring the retrieval process based on the complexity of the questions. This approach improves alignment between user intent and retrieved content by considering the model’s responses alongside RAG effects. However, measuring the effectiveness of retrieved information on model outputs remains challenging, especially in complex and open-domain scenarios.
Given these challenges, using GPT-4-based classification labels for questions to generate data for Llama-2 7B evaluations is not reasonable, as it does not consider the knowledge boundaries of Llama-2 7B. In our analysis, we find that performance varies significantly among LLMs of different sizes, indicating the need to focus on specific models for a more accurate simulation of their knowledge boundaries, rather than just distilling larger models for labels. Additionally, generating soft labels from queries and LLM predictions often requires gold answers for evaluation, posing challenges for questions without definitive answers. For questions with gold answers, we sample multiple LLM responses to compute confidence levels. In contrast, for questions lacking gold answers, we evaluate the entropy of phrase distributions from the model’s responses to create a certainty distribution. Based on these evaluation indicators, we categorize RAG’s impact on LLMs as beneficial, neutral, or harmful based on these evaluation indicators. Our findings show that as the LLM’s confidence or certainty increases, the ratio of neutral and harmful queries also rises, while beneficial queries and performance improvements decline. This motivates us to reduce retrieval requests leading to neutral or harmful outcomes, optimizing time and computational efficiency while maximizing beneficial retrievals to enhance question answering (QA) performance.
In this work, we propose methods to evaluate the known and unknown aspects of LLMs based on indices of confidence and certainty derived from sampled responses. By establishing different thresholds for confidence and certainty, we can effectively distinguish between beneficial questions and those deemed neutral or harmful, thereby enabling the generation of data with soft labels. The data is used to fine-tune the Knowledge Boundary Model (KBM) to determine if the LLM considers the question unknown, thus indicating whether retrieval is necessary. This assessment informs decisions on whether RAG enhancement is necessary, leading to better performance and decreased retrieval costs. We evaluate several English and Chinese QA datasets and demonstrate that KBM effectively detects knowledge mastery in LLMs through end-to-end results. Our confidence-based approach significantly reduces retrieval ratios while achieving performance close to ALL RAG, with a 43.17% reduction in retrieval on WebQA and a performance drop of only 1.7%. Similarly, the certainty-based approach maintains performance near ALL RAG (with a 0.42% decrease) and reduces the search ratio by approximately 10% on QA tasks. Further analysis explores KBM’s effectiveness in open domains across three dimensions: dynamic knowledge, long-tail static knowledge, and multi-hop problems. Finally, we discuss the impact of incorporating RAG on the knowledge boundaries of LLMs and potential solutions based on KBM to address these challenges.
Figure 1: Illustration of the impact of RAG on LLM
performance.
Figure 2: Illustration of the ratio of different LLM output MASK.
2 Related work
Large Language Model Knowledge Exploration. The exploration of knowledge boundaries in LLMs attracts significant attention. Kadavath et al. (2022) examines the self-evaluation capabilities of LLMs, showing that larger models enhance their calibration by initially proposing answers and then evaluating their validity.However, their model still face challenges in accurately identifying errors. Ren et al. (2023) studies LLMs’ perception of factual knowledge boundaries and finds that they often display blind confidence in their abilities. Using retrieval-augmented questioning, the study shows that retrieval can enhance judgment capabilities but highlights a dependence on the quality of supporting documents. Yin et al. (2023) focuses on self-awareness, demonstrating that while LLMs can identify some unanswered questions, substantial discrepancies still exist, affecting their uncertainty detection. Chen et al. (2024) introduces COKE, an unsupervised method for teaching models to articulate their knowledge limits through internal signals, yielding improved outcomes across various datasets. Kang et al. (2024) points out that LLMs often default to examples in training data when facing unfamiliar queries. This highlights the need for better factuality controls to reduce instances of hallucination. Li et al. (2024b) explores hallucinations related to insufficient prompt context, showing that models frequently fail to recognize inadequate information. They propose uncertainty-sensitive tuning to improve response reliability. Based on the above research work, our research introduces the KBM. This model aims to enhance the ability of LLMs to determine when and what external knowledge to retrieve for open-domain QA, revealing the key dynamics between internal and external knowledge utilization.
Retrieval-Augmented Generation. RAG enhances LLMs by integrating retrieved text passages, significantly improving performance in knowledge-intensive tasks. A key focus is optimizing the timing and strategy of retrieval. Asai et al. (2023) introduce SELF-RAG, a method that trains LLMs to retrieve information, generate content, and evaluate their outputs using reflection tokens. This method enables the customization of model behavior, demonstrating significant performance improvements over standard RAG approaches. Jeong et al. (2024) proposes Adaptive-RAG, which adjusts query handling based on complexity. The model employ different strategies, including non-retrieval and multi-step approaches,to enhance the relevance of responses to diverse queries. Wu et al. (2024) explores how LLMs process erroneous retrieved content. By creating a dataset to assess model responses to incorrect information, the study reveals insights into how models correct their outputs or may perpetuate errors. Cuconasu et al. (2024) conducts a comprehensive study on the retriever’s function in RAG. Key findings reveal that the positioning of relevant documents affects model performance, and unexpectedly, noisy documents can enhance accuracy when placed strategically. These studies demonstrate the ongoing development of RAG methodologies, laying the groundwork for our investigation into new strategies for effective retrieval during generation.
3 PRELIMINARIES: LLM Knowledge Boundaries and RAG Analysis
This section examines RAG’s impact on LLM performance. We first assess LLMs with varying parameter counts on QA tasks , revealing differences in knowledge boundaries related to parameter size. Next, we evaluate how each LLM utilizes textual information retrieved by naive RAG and investigate the effects of MASK perturbation to understand their reliance on RAG. We categorize questions into three types based on RAG’s influence on LLM outputs. Additionally, we propose using confidence based on LLM accuracy and the certainty of output words as indicators, motivating further study into simulating knowledge boundaries with these evaluation methods.
3.1 How does RAG affect the accuracy and uncertainty of LLM response?
We assess the impact of RAG on the performance of LLMs with varying parameter sizes. Our findings indicate that while LLMs exhibit different abilities to answer questions and possess distinct knowledge boundaries, their capacity to utilize retrieved text information remains largely consistent. Our analysis employs several configurations: LLM Only: This configuration generates responses directly from the LLM. ALL RAG: We enhance the Naive RAG approach by concatenating the top ten blocks retrieved from Google as contextual information. MASK RAG: This variant replaces the gold answers present in RAG with MASK, supplying these modified data as context for the LLM.
@@ 3.1 How does RAG affect the accuracy and uncertainty of LLM response?
As illustrated in Figure 2, we focus on the Qwen1.5 models (4B, 7B, 14B, 32B) alongside the Qwen- 2.0 72B model. The evaluation datasets utilized for assessing short question answering and reading comprehension tasks include WebQA (Chang et al., 2022), SogouQA [^1] https://github.com/sherlcok314159/ChineseMRC-Data, and SQuAD1.5-zh [^2] https://github.com/pluto-junzeng/ChineseSquad. Our results demonstrate that LLMs improve their performance with increasing parameter sizes across the three datasets, exhibiting a gradual differentiation in their QA capabilities. Notably, when employing Naive RAG, all models, particularly those with 14B-72B parameters, display a strong and consistent ability to exploit contextual information. We find that RAG has a more pronounced effect on smaller LLMs, although the upper limit of performance improvement remains similar across model sizes. For instance, in the WebQA dataset, the accuracy difference between the 7B and 72B models when using RAG is only 3.05%. In contrast, without RAG, this difference escalates to 28.87%. Interestingly, the use of MASK RAG appears to diminish the advantages associated with RAG, potentially leading to performance degradation, especially on simpler datasets. The presence of noisy information adversely impacts smaller models to a greater extent.
@@
These findings suggest that different LLMs possess varying knowledge boundaries in question answering and demonstrate distinct retrieval strategies. Although all models exhibit strong capabilities in leveraging context, their resilience to noise interference varies significantly.
(a) WebQA Acc. LLM Only
(b) WebQA Acc. ALL RAG
(c) WebQA Acc. MASK RAG
(d) WebQA Cer. LLM Only
(e) WebQA Cer. ALL RAG
(f) WebQA Cer. MASK RAG
Figure 3: Predictions for three categories: Beneficial, Neutral, and Harmful on the WebQA test set. The top panel shows accuracy distribution; the bottom panel shows certainty distribution. The left section, LLM Only, shows scores for the three categories. The middle section, ALL RAG, presents scores from the RAG approach. The right section displays scores for MASK RAG.
3.2 Three dimensions to divide the boundaries of RAG
@@
To more precisely evaluate the impact of RAG on LLM generation, we categorize its effects into three aspects:
  • Beneficial: RAG effectively solves this problem.
  • Neutral: RAG does not impact the LLM’s effectiveness.
  • Harmful: RAG compromises the LLM’s effectiveness.
@@
To isolate the influences of other modules, we utilize a simplified RAG pipeline for our analysis. Specifically, we leverage Google Open Search to retrieve the top 10 blocks, which are then provided to the LLM. Acknowledging that a single model response may lack robustness or representativeness, we implement a confidence-based categorization method. Inspired by Kadavath et al. (2022), we generate $I=30$ answer samples at $T=1$. For a given question $Q$, if 20 of the answers sampled by the model are correct and 10 are wrong, unlike Kadavath et al. (2022) which contains 20 copies of $(Q,M=1)$ data points and 10 copies of $(Q,M=0)$ data points. Instead, we construct a confidence of knowledge for this question based on our samples, resulting in a single data point $(Q,M_{pred}=\frac{20}{30})$. This method aims to accurately represent the model’s understanding and misconceptions while significantly reducing the size of the training and test datasets by a factor of $I$. Consequently, we approximate the model’s soft labels for knowledge using hard labels derived from a diverse set of QA data points.
@@
However, this method becomes challenging in the absence of gold-standard answers. To address this, we simulate the effect based on the uncertainty of the generated responses. We compute the word/phrase entropy distribution of the LLM based on the 30 generated answers to assess the model’s certainty. Let $k$ represent the number of distinct answer types, denoted as $K_{1},K_{2},...,K_{k}$. The probability of each answer type occurring is represented as $P_{1},P_{2},...,P_{k}$, satisfying the normalization condition: $\sum_{i=1}^{k}P_{i}=1$. Using this probability distribution, we quantify certainty through entropy with the following formula:
$C(Q)=1-U_{PE}(Q)=1+\sum_{i=1}^{k}P_{i},{\mathrm{log}}\,P_{i}.$ (1)
where $U_{PE}\in[0,1]$ represents the entropy value, which quantifies the model’s uncertainty or confidence in its answers.
We classify the effects of RAG based on the three categories and two indicators as follows: an effect is deemed Beneficial if the indicator increases after incorporating RAG compared to the LLM Only scenario; Neutral if the indicator remains unchanged and aligns with the value from the LLM Only condition; and Harmful if the indicator decreases following the addition of RAG, falling below the value observed in the LLM Only scenario. Using these categories, we analyze the class distribution of LLMs across accuracy intervals on the WebQA and SogouQA datasets, as shown in Figure 3. The figure indicates that the [0-0.8) interval has the highest proportion of beneficial cases, while neutral and harmful cases peak in the [0.8-1.0] range. This suggests a cost-benefit relationship: higher accuracy reflects greater confidence in answers, reducing the advantages of RAG, especially in the [0.8-1.0] range, where harmful cases also concentrate.
When an LLM displays high confidence, we infer mastery of the solution; low confidence implies a lack of understanding. This assessment combines human annotations, model predictions, and evaluation methods. The uncertainty metric highlights components such as Aleatoric and Model Uncertainty. Low uncertainty corresponds to high confidence, while high uncertainty indicates doubt. Thus, determining mastery can significantly lower retrieval needs, as illustrated in the blue area of the figure. For instance, with mastery defined as accuracy exceeding 0.8, a 7B model on WebQA reduces retrieval by 55.47% and enhances performance by 0.92%. A 72B model achieves an 80.95% reduction in retrieval and a 2.52% performance gain. If we set a confidence threshold of under 0.2, the 7B model reduces retrieval by 41.45%, but this results in a 4.15% performance decline compared to AllRAG. The 72B model achieves a 61.56% reduction with a 1.56% drop in performance. We observe a Pearson correlation coefficient of 0.64 between confidence and certainty, indicating a positive relationship. However, certainty-based metrics tend to be more inclusive.
Overall, the thresholds for mastery and confidence depend on the specific dataset and model. While accuracy aligns with human intent and reflects model capabilities, its reliability is affected by annotation quality and evaluation variability. Uncertainty metrics also have limitations, particularly concerning model overconfidence, which can degrade performance. These aspects will be further examined in the following sections.
3.3 The dependence of LLM on RAG
In this section, we explore the reliance on RAG from the perspective of LLM uncertainty. We find that incorporating RAG significantly enhances LLM certainty, which is evident in the greater consistency of the generated answers. This observation raises a critical question: does the model primarily copy answers from the RAG input to reduce uncertainty, or does it absorb knowledge from the provided material? To address this question, we conduct an experiment with scenarios where the model outputs exhibit a certainty of $1$, indicating high confidence in predictions. As shown in Figure 2, we measure the frequency of outputs corresponding to MASK under MASK RAG. If the LLM mainly copies answers from the RAG input, we expect to observe a high proportion of MASK outputs; this would imply that the model retrieves and reproduces answers rather than inferring them from context. Our findings are consistent across WebQA, SogouQA and SQuAD-zh datasets. Notably, the tendency to copy answers decreases with an increase in the number of model parameters, leading to a lower proportion of MASK outputs. This suggests that larger models are more adept at leveraging non-answer content to derive predictions, rather than relying on direct answer extraction. Such reliance demands more from RAG and may negatively affect LLM performance.

### v2

# KBM: Delineating Knowledge Boundary for Adaptive Retrieval in Large Language Models

@@ Abstract
Large Language Models (LLMs) often struggle with dynamically changing knowledge and handling unknown static information. Retrieval-Augmented Generation (RAG) is employed to tackle these challenges and has a significant impact on improving LLM performance. In fact, we find that not all questions need to trigger RAG. By retrieving parts of knowledge unknown to the LLM and allowing the LLM to answer the rest, we can effectively reduce both time and computational costs. In our work, we propose a Knowledge Boundary Model (KBM) to express the known/unknown of a given question, and to determine whether a RAG needs to be triggered. Experiments conducted on 11 English and Chinese datasets illustrate that the KBM effectively delineates the knowledge boundary, significantly decreasing the proportion of retrievals required for optimal end-to-end performance. Furthermore, we evaluate the effectiveness of KBM in three complex scenarios: dynamic knowledge, long-tail static knowledge, and multi-hop problems, as well as its functionality as an external LLM plug-in.
KBM: Delineating Knowledge Boundary for Adaptive Retrieval in
Large Language Models
Zhen Zhang1, Xinyu Wang2thanks:   Corresponding author, Yong Jiang2∗, Zile Qiao2, Zhuo Chen3, Guangyu Li3, Feiteng Mu, Mengting Hu1∗, Pengjun Xie2, Fei Huang2 1College of Software, Nankai University 2Tongyi Lab, Alibaba Group 3School of Information Science and Technology, ShanghaiTech University 3Shanghai Engineering Research Center of Intelligent Vision and Imaging zhzhen23@gmail.com
@@ 1 Introduction
As Large Language Models (LLMs) evolve, their real-world applications expand, yet they often struggle with dynamically changing and unknown static knowledge, leading to inaccuracies or hallucinations (Rawte et al., 2023). Retrieval-Augmented Generation (RAG) effectively addresses challenges by retrieving relevant external information in real time, enhancing LLMs’ accuracy. However, it also incurs costs, such as increased retrieval requests and longer response times, leading to the crucial question: When is retrieval truly necessary?
Figure 1: Illustration of the impact of different RAG triggering methods on LLM. Beneficial: RAG effectively solves this question. Neutral: RAG has no effect on the LLM. Harmful: RAG reduces the LLM’s effectiveness.
A natural criterion for determining the need for retrieval is the confidence level of the LLM. Specifically, the known knowledge of the model does not require retrieval, while uncertain parts can benefit from this process. As illustrated in Figure 1 (top), using RAG for all questions (All RAG) can result in an overall impact that is neutral or even harmful, as the increased retrieval calls may not justify the benefits. In contrast, Figure 1 (bottom) demonstrates that when focusing solely on the uncertain part through RAG, the benefits are substantial, while reducing the frequency of unnecessary retrieval calls. The key to implementing this solution is how to enable LLM express confidence. A simple and effective solution is to use prompts. Unfortunately, over-/under-confidence will affect the expression of confidence Xiong et al. (2024).
@@
Previous studies on whether RAG is required for LLM can be divided mainly into two approaches. The first focuses on the question itself, with methods like Self-RAG (Asai et al., 2023) instructing models such as GPT-4 (Achiam et al., 2023) to assess whether retrieving external documents (e.g., Wikipedia) can produce better responses. Although this approach can identify questions that require real-time information, it remains model-agnostic and struggles to determine whether an LLM has mastered specific knowledge. The second approach evaluates both questions and model responses to determine if an LLM can answer a question, generating data by sampling multiple model responses and using manual labels for evaluation Ren et al. (2024); Yin et al. (2023); Kadavath et al. (2022). However, this method is labor intensive and relies heavily on manual labeling, which can create biases and lead to increased training costs. Additionally, the LLM’s "unknown" expression has not been integrated with RAG triggering to assess its impact on end-to-end performance and retrieval ratio.
In this work, we propose the Knowledge Boundary Model (KBM) and two solutions for SFT based on accuracy and certainty in generating soft labels. The first method evaluates sampled QA accuracy and sets a threshold to classify questions as known or unknown. The second approach is based on certainty, focusing on modeling the entropy values of multiple sampled responses to establish thresholds, distinguishing between known and unknown responses through consistency, without requiring a golden answer. By defining different accuracy and certainty thresholds, we can generate data with soft labels, which is then used to fine-tune the KBM to determine whether the LLM considers a question unknown, thus indicating if retrieval is necessary.
We evaluate 11 English and Chinese QA datasets, demonstrating that KBM effectively measures knowledge mastery in LLMs. On average in these datasets, the certainty-based approach also reduces the number of retrievals by 13. 5% compared with ALL RAG, resulting in a slightly higher performance of 0.1%. Similarly, the accuracy-based approach reduces retrievals by 32.5% on average, with only a 0.39% performance drop. Further analysis confirms the effectiveness of KBM in open domains, including dynamic knowledge, long-tail static knowledge, and multi-hop scenarios.
Our contributions are summarized as follows:
  • To our knowledge, KBM is the first study to reduce the RAG trigger ratio while maintaining LLM performance, thereby enhancing RAG’s efficiency in QA tasks and reducing costs.
  • For the technique contribution, we propose two methods for generating soft labels based on accuracy and certainty, allowing LLMs to express "unknown" or "known".
  • KBM is validated on 11 datasets, demonstrating comparable effectiveness with All RAG and a reduced retrieval ratio, and it performs well in three complex scenarios.
2 Preliminaries: LLM Knowledge Boundary and RAG Analysis
This section examines the impact of RAG on LLM performance. We evaluate the knowledge boundaries of LLMs with different parameter sizes and analyze RAG’s effects on QA tasks, categorizing questions into three types based on its influence on LLM responses. Lastly, we introduce sampling methods based on LLM accuracy and certainty to further simulate knowledge boundary.
2.1 How does RAG Affect the Accuracy of LLM Response?
We evaluate how RAG affects the performance of LLMs with different parameter sizes. Our results show that while LLMs vary in their QA abilities and knowledge limits, their use of retrieved information is fairly consistent. We use three configurations: LLM Only generates responses from the LLM alone; ALL RAG enhances Naive RAG by adding the top 10 blocks retrieved from Google as context; and MASK RAG substitutes the correct answers in RAG with MASK, providing this altered data as context for the LLM.
@@ 2.1 How does RAG Affect the Accuracy of LLM Response?
We focus on the Qwen1.5 models (4B, 7B, 14B, 32B) Bai et al. (2023) alongside the Qwen2.0 72B model Yang et al. (2024), using evaluation datasets for short question answering and reading comprehension tasks, including WebQA (Chang et al., 2022), SogouQA [^1] https://github.com/sherlcok314159/ChineseMRC-Data, and SQuAD1.5-zh [^2] https://github.com/pluto-junzeng/ChineseSquad. Our results show that LLM performance improves with larger parameter sizes across all three datasets, leading to a gradual differentiation in their QA capabilities. Notably, with Naive RAG, all models, especially those with 14B-72B parameters, effectively utilize contextual information. RAG has a more significant impact on smaller LLMs, but the maximum performance gain is similar across different model sizes. For example, on the WebQA dataset, the accuracy difference between the 7B and 72B models using RAG is just 3.05%, compared to 28.87% without RAG. Interestingly, MASK RAG seems to reduce the benefits of RAG, potentially harming performance on simpler datasets, as noisy information negatively affects smaller models more. For detailed results, see Appendix §A Figure 9.
@@
These findings suggest that different LLMs possess varying knowledge boundary in question answering and demonstrate distinct retrieval strategies. Although all models exhibit strong capabilities in leveraging context, they show varying degrees of resilience to noise interference.
2.2 Three Impacts of RAG on LLM
@@
To accurately assess the impact of RAG on LLMs, we categorize its effects into three aspects:
  • Beneficial: RAG effectively solves this question.
  • Neutral: RAG has no effect on the LLM.
  • Harmful: RAG reduces the LLM’s effectiveness.
@@
To isolate the impact of other modules, we use a simplified RAG pipeline for analysis, retrieving the top 10 blocks from Google Open Search to input into the LLM. To improve robustness, we sample each question 30 times. Inspired by Kadavath et al. (2022), we generate $I=30$ answer samples at temperature $T=1$. For a given question $Q$, if the model samples 20 correct answers and 10 incorrect ones, we construct an average accuracy of knowledge for that question based on these samples, resulting in a single data point $(Q,M_{pred}=\frac{20}{30})$. Unlike the approach in Kadavath et al. (2022), which includes 20 copies of (Q, M=1) and 10 copies of (Q, M=0) our method accurately represents the model’s understanding and misconceptions while significantly reducing the size of the training and test datasets by a factor of $I$. Thus, we approximate the model’s soft labels for knowledge using hard labels derived from a diverse set of QA data points.
@@
However, this method becomes challenging in the absence of gold-standard answers. To address this issue, we simulate the effect based on the certainty of the generated responses. Specifically, we compute the entropy distribution of words and phrases from the 30 generated answers to establish the model’s level of certainty. Let $k$ represent the number of distinct answer types, denoted as $K_{1},K_{2},...,K_{k}$. The probability of each answer type occurring is represented as $P_{1},P_{2},...,P_{k}$, satisfying the normalization condition: $\sum_{i=1}^{k}P_{i}=1$. Using this probability distribution, we quantify the model’s certainty through entropy. First, we define entropy $H$ as:
(a) Category distribution based on Accuracy interval.
(b) Category distribution based on Certainty interval.
Figure 2: Category distribution across various indicators: (a) Accuracy and (b) Certainty intervals.
@@
$H=-\sum_{i=1}^{k}P_{i}\log_{2}(P_{i}),$ (1)
and the maximum possible entropy $H_{max}=\log_{2}(k).$ The model’s certainty $C(Q)$ is calculated by normalizing the entropy:
$C(Q)=1-\frac{H}{H_{{max}}},$ (2)
where ${H}/H_{max}\in[0,1]$ is the normalized entropy.
We classify the effects of RAG based on the three categories and two indicators as follows: an effect is deemed Beneficial if the indicator increases after incorporating RAG compared to the LLM Only scenario; Neutral if the indicator remains unchanged and aligns with the value from the LLM Only condition; and Harmful if the indicator decreases following the addition of RAG, falling below the value observed in the LLM Only scenario. Using these categories, we analyze the class distribution of LLMs across accuracy intervals on the WebQA dataset, as shown in Figure 2(a). The figure indicates that the [0-0.8) interval has the highest proportion of Beneficial cases, while Neutral and Harmful cases peak in the [0.8-1.0] range. This suggests a cost-benefit relationship: higher accuracy reflects greater confidence in answers, reducing the advantages of RAG, especially in the [0.8-1.0] range, where harmful cases also concentrate. High certainty reflects high confidence, while low certainty indicates doubt. The certainty metric considers factors such as chance and model uncertainty Hu et al. (2023). We find a Pearson correlation coefficient of $0.64$ between accuracy and certainty, as shown in Figure 2. Certainty also has a similar category distribution to accuracy, indicating a positive correlation.
In the high accuracy range (above $0.8$), we first note that many neutral examples unnecessarily trigger RAG. Additionally, harmful examples can lead to a sharp accuracy drop, potentially from $1.0$ to $0.0$, while beneficial examples typically increase accuracy from 0.8 to a maximum of $1.0$. This leads us to propose a threshold to differentiate between known and unknown data. Data below this threshold is marked as unknown and triggers RAG, while data above it is considered known, allowing the LLM to respond independently.